A normal hydrogen atom consist of a proton nucleus and oribital electron , each carrying a charge (e).assuming that the electron orbit is circular and the seperation between the particles is 5.3×10^-11m.find

(I) the force of electrical attraction between the particles

(II) the electron oribital speed


1
Expert's answer
2022-02-02T05:14:20-0500

Given:

"r=5.3×10^{-11}\:\rm m"

"e=1.6×10^{-19}\:\rm C"

"m=9.1×10^{-31}\:\rm kg"


(I) the force of electrical attraction between the particles

"F=k\frac{e^2}{r^2}=9*10^9\frac{(1.6*10^{-19})^2}{(5.3×10^{-11})^2}=8.2*10^{-6}\: \rm N"

(II) the electron oribital speed

"F=\frac{mv^2}{r}"

"v=\sqrt{\frac{Fr}{m}}=\sqrt{\frac{8.2*10^{-6}*5.3×10^{-11}}{9.1×10^{-31}}}=2.2*10^7\:\rm m/s"